The tangents at the points P(ap2 , 2ap) and Q(aq2 , 2aq) on the parabola y2= 4ax intersect at the point R. Given that the tangent at P is perpendicular to the chord OQ, where O is the origin, find the equation of the locus of R as p varies.


1
Expert's answer
2021-08-17T10:37:31-0400

"P(ap^2,2ap), Q(aq^2,2aq)\\
y^2=4ax"

The tangents is

"y\cdot y_0=2a(x+x_0)"

PR:

"y\cdot 2ap=2a(x+ap^2)\\
x-py+ap^2=0"

QR:

"y\cdot 2aq=2a(x+aq^2)\\
x-qy+aq^2=0"

R=PR"\cap" QR

"x-py+ap^2=0\\
x-qy+aq^2=0\\
y=\frac{a(p^2-q^2)}{q-p}=a(p+q)\\
x=aqp\\
R(aqp,a(p+q))"

OQ:

"\frac{x-0}{aq^2-0}=\frac{y-0}{2aq-0}\\
y=\frac{2}{q}x\\
k_{OQ}=\frac{2}{q}"

OQ"\bot" RP

"k_{RP}=-\frac{1}{k_{OQ}}=-\frac{q}{2}"

RP:

"y=\frac{1}{p}x+ap\\
k_{RP}=\frac{1}{p}\\
\frac{1}{p}=-\frac{q}{2}\\
pq=-2"


multiply by a

"apq=-2a\\
x=-2a\\
x+2a=0"